The mean weight gain during pregnancy is 30 pounds with a standard deviation of 12.9 pounds. If an obstetrician obtains a random sample of 35 patients, what is the probability that the mean weight gain during pregnancy from that sample is at least 34.5 pounds?

That is, with n = 35, what is P(x≥34.5)?
Answer #1

Solution :

Given that,

mean = = 30

standard deviation = =12.9

n=35

= =30

= / n = 12.9 / 30 = 2.3552

P(x≥34.5) = 1 - P( <34.5 )

= 1 - P[( - ) / < (34.5-30) /2.3552 ]

= 1 - P(z <1.91 )

Using z table

= 1 - 0.9719

= 0.0281

probability= 0.0281